A 25-foot ladder leaning against a building touches the wall 18 feet above the ground. How far from the building is the bottom of the ladder? Help please. Thank you
Answer by checkley75(3666)

THIS PROBLEM IS A RIGHT TRIANGLE WITH THE HYPOTENUSE OF 25 & ONE SIDE IS 18. THEN:
25^2=16^2+X^2
625=256+X^2
X^2=625-256
X^2=369
X=SQRT369
X=19.2 FEET FOR THE DISTANCE FROM THE BOTTOM OF THE LADDER TO THE SIDE OF THE BUILDING.